OpenStudy (naomi):

If y = 6x − 3, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? {(0, 3), (1, −9), (2, −15)} {(0, −3), (2, 15), (4, 21)} {(0, −3), (2, 9), (4, 21)} {(0, 3), (−2, 15), (4, 21)}

pooja195 (pooja195):

Hi there! Welcome to OpenStudy :) Let's try The first choice, we are going to plug in the x and y values into the given equation

pooja195 (pooja195):

3=6(0)-3 -9 = 6(1) − 3 -15= 6(2) − 3 Can you tell me if these ordered pairs make the equation true? (these are from choice A)

OpenStudy (naomi):

?

pooja195 (pooja195):

3=6(0)-3 Is this true? 6*0-3=3?

OpenStudy (naomi):

it would be -3

pooja195 (pooja195):

Good so we know that A. is not the answer

pooja195 (pooja195):

Let's try B {(0, −3), (2, 15), (4, 21)} -3=6(0)-3 15=6(2)-3 21=6(4)-3 Are these true?

OpenStudy (naomi):

the second one isnt

pooja195 (pooja195):

Good so we know B is not our answer, lets try C {(0, −3), (2, 9), (4, 21)} We already know the (0,-3) and (4,21) are right so we only need to try (2,9) 9=6(2)-3 is this true?

OpenStudy (naomi):

idk

pooja195 (pooja195):

Does 6*2-3=9?

OpenStudy (naomi):

._.

pooja195 (pooja195):

?

OpenStudy (naomi):

yes

pooja195 (pooja195):

Good so which of the choices is out answer?

OpenStudy (naomi):

c?

pooja195 (pooja195):

Yep well done :)

OpenStudy (naomi):

cool thank you and you changed your picture

pooja195 (pooja195):

Yep and wanted a new look :P ^_^

OpenStudy (naomi):

thx